OpenStudy (jazzyfa30):

please help Exponential and Logarithmic Equations solve each equation. Give an exact solution. 8^(x+2)=16

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

good this is we dont need to mess with logs for doing this q

ganeshie8 (ganeshie8):

il give u a hint : 8 = 2^3

ganeshie8 (ganeshie8):

another hint : 16 = 2^4

ganeshie8 (ganeshie8):

\(\large 8^{x+2}=16\)

ganeshie8 (ganeshie8):

write 8 and 16 as powers of 2

ganeshie8 (ganeshie8):

\(\large 2^{3(x+2)}=2^4\)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (jazzyfa30):

yes

ganeshie8 (ganeshie8):

\(\large 2^{3(x+2)}=2^4\) since bases are equal on both sides, we can equate the exponents : \(\large 3(x+2) = 4\)

ganeshie8 (ganeshie8):

solve \(x\)

OpenStudy (jazzyfa30):

x=(-2/3)

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)